# The Peripatetic Hater: Predicting Movement Among Hate Subreddits

### v1

@@ Abstract
Many online hate groups exist to disparage others based on race, gender identity, sex, or other characteristics. The accessibility of these communities allows users to join multiple types of hate groups (e.g., a racist community and misogynistic community), which calls into question whether these peripatetic users could be further radicalized compared to users that stay in one type of hate group. However, little is known about the dynamics of joining multiple types of hate groups, nor the effect of these groups on peripatetic users. In this paper, we develop a new method to classify hate subreddits, and the identities they disparage, which we use to better understand how users become peripatetic (join different types of hate subreddits). The hate classification technique utilizes human-validated LLMs to extract the protected identities attacked, if any, across 168 subreddits. We then cluster identity-attacking subreddits to discover three broad categories of hate: racist, anti-LGBTQ, and misogynistic. We show that becoming active in a user’s first hate subreddit can cause them to become active in additional hate subreddits of a different category. We also find that users who join additional hate subreddits, especially of a different category, become more active in hate subreddits as a whole and develop a wider hate group lexicon. We are therefore motivated to train an AI model that we find usefully predicts the hate categories users will become active in based on post text read and written. The accuracy of this model may be partly driven by peripatetic users often using the language of hate subreddits they eventually join. Overall, these results highlight the unique risks associated with hate communities on a social media platform, as discussion of alternative targets of hate may lead users to target more protected identities.
@@ Introduction
Recently, there has been extensive work documenting online hate communities and the impact that they have on the users of social media platforms (Schmitz et al. 2024; Russo et al. 2023). Such research is motivated by the combination of substantial increases in hate crimes (of Investigation 2021), and the observation that radicalized individuals attribute their beliefs to online sources (Jensen et al. 2018). Much of the existing literature in this domain provides either case studies of individual communities (Chandrasekharan et al. 2017; Schmitz et al. 2024; Russo et al. 2023) or broad analyses of groups centered around a common theme, such as the manosphere (Ribeiro et al. 2021). Additionally, although prior work observes strong user overlap between anti-feminism and alt-right communities (Mamié, Horta Ribeiro, and West 2021), such research does not show how the core themes discussed in these communities are influenced by the users active in both communities.
Here, we extend the existing research corpus by curating a comprehensive dataset of 120 hate subreddits and using LLMs and clustering techniques to group them by ideology (racist, misogynistic, and anti-LGBTQ). In these subreddits, we analyze peripatetic users, who participate in multiple categories of hate subreddits, attempting to characterize them and understand the impact that they have on Reddit’s ecosystem of hate. Our research questions are as follows:
@@ 4. RQ4
Our findings help answer each respective research question: (RQ1) Using a matched-pair causal model, we find that participation in a given hate subreddit is associated with joining other hate subreddits at a rate roughly seven times higher than that of control users. (RQ2) Using BERTopic (Grootendorst 2022) and lexicons of ingroup language, we find that peripatetic users introduce into their initial subreddit topics of discussion from the categories in which they eventually participate. For example, a user who initially posts only in a misogynistic subreddit, but subsequently also posts in a racist subreddit, is more likely to introduce racist topics into the misogynistic subreddit. (RQ3) We find users that join additional hate subreddits become more active in hate subreddits as a whole and use lexicons associated with these newly joined subreddits. Both results are more pronounced when users join hate subreddits of a different category then the first subreddit they joined. (RQ4) Finally, we develop a deep-learning model to predict the subreddit categories these peripatetic users will join. This task is challenging, as we are differentiating among users all of whom already exhibit signs of hostility. We find that the language of users who eventually become peripatetic is strongly associated with the subreddit categories that they subsequently join.
Overall, our results offer a novel perspective on how hate communities may converge and become more “general” in an online environment as a result of peripatetic user activity. All data and code are in the following repository https://github.com/dan-hickey1/peripatetic-hater.
@@
Misogynist. Extensive work has investigated the various identities targeted by online hate communities. One of the most widely studied groups is the manosphere, a collection of communities centered around misogyny and anti-feminism (Marwick and Caplan 2018; Habib, Srinivasan, and Nithyanand 2022). Despite sharing this common theme, there is a wide diversity of ideologies within the manosphere. Incels, for example, believe they are incapable of finding a sexual or romantic partner and blame women and society for such misfortune (Daly and Reed 2022). This is in contrast with pick-up artists, who discuss strategies for seducing women, and “men going their own way” (MGTOW), who voluntarily decide to be celibate as they believe they should distance themselves from women (Lin 2017). A recent analysis shows MGTOW and incel communities are growing in popularity and are more toxic than other manosphere communities, which are decreasing in popularity (Ribeiro et al. 2021). Critically, members of incel communities have engaged in mass shooting events (Barcellona 2022).
Anti-LGBTQ. Anti-LGBTQ communities are also prevalent on the internet, including trans-exclusionary radical feminists (TERFs), an anti-transgender group. TERFs generally subscribe to a binary view of gender that is aligned with biological sex (disregarding intersex individuals), and see transgender people as a threat to women’s spaces (Jones and Slater 2020). Scholars have previously demonstrated flaws in such discourse and the harm it produces (Williams 2020; Hotine 2021). TERF communities have been present on Reddit in the past, with r/GenderCritical being the most prominent (Tiffany 2020). Previous work has found these communities to be more toxic than mainstream feminist communities (Balci et al. 2023), and models have been built to automatically identify the discourse employed by them (Lu and Jurgens 2022).
Racist. Racist groups have been highly prevalent on the internet, and on Reddit in particular (Chandrasekharan et al. 2017), as have a uniquely online group known as the alt-right, an extreme political movement that gained popularity through its utilization of the unique characteristics of social media (Winter 2019). The alt-right believe that the cultural identity of white individuals is at risk, with movements for social justice and political correctness being seen as principal threats (Center 2023). Alt-right online communities have radicalized users in part through the use of memes (Dafaure 2020), with one analysis showing that communities linked to the alt-right have a strong influence on the spread of memes in the larger web ecosystem (Zannettou et al. 2018). The misogynist movement “GamerGate” helped the alt-right movement gain adherents (Bezio 2018), demonstrating a relationship between seemingly distinct types of hate.
@@
Radicalization is largely explicable in terms of three non-exclusive effects: (1) uncertainty reduction (Hogg, Kruglanski, and van den Bos 2013), (2) significance quest (Kruglanski et al. 2014), and (3) the need to be popular (Siegel et al. 2011). There is also substantial research into the micro-scale processes that drive people to radicalization. From a community-level perspective, scholars have studied whether certain communities are “gateways” to more extreme communities, such as the connection between the intellectual dark web and the alt-right (Ribeiro et al. 2020), or between the manosphere and the alt-right (Ribeiro et al. 2020; Barcellona 2022). Gateway communities to political and conspiracy subreddits have been automatically identified using a network analysis approach (Rollo et al. 2022). Additional research on conspiracy communities demonstrates how users who increase their engagement in specific conspiracy theory communities over time also increase their participation in less specialized conspiracy theory communities (Phadke, Samory, and Mitra 2022). Other studies measure traits indicative of user radicalization or extremism that are adopted by members of certain hate groups, finding that such traits increase once users participate in hate subreddits (Habib, Srinivasan, and Nithyanand 2022; Schmitz et al. 2024). While recommendation algorithms are often blamed for radicalizing users, recent evidence shows that users who consume extreme content on YouTube are driven to do so by external sources, and they already have high racial and gender resentment before engaging in such content (Chen et al. 2023).
@@
There is a growing body of literature devoted to understanding how users participate in multiple communities on the same platform. One theme in this research is the attempt to understand inter-community conflict, where members active in one community make posts in other communities to attack them (Kumar et al. 2018; Efstratiou et al. 2023; Datta and Adar 2019). Another paper speculates that mass migrations of users among communities on Reddit may disrupt the linguistic evolution of destination communities, which could influence their likelihood of being banned (Habib et al. 2021). More generally, Rollo et al. (Rollo et al. 2022) devise a model of “attention flow” to quantitatively describe how users on Reddit shift their focus among subreddits over time. Scholars have also built models to predict whether users will participate in particular online communities (Jin, Lee, and Cheung 2010; Joyce and Kraut 2006; Debaere, Coussement, and De Ruyck 2018), with one study finding that the social support is predictive of users remaining in an online health community (Wang et al. 2017).
The present study builds on previous research by examining a multitude of hate communities while assessing the connections among all of them. In addition, we describe the tendency of users to join specific additional hate communities after initially joining a given hate community and measure the user behavior that predates such additional participation. Finally, we build a model to predict which new hate subreddits users will join.
@@ Methods
To properly understand how different categories of hate subreddits influence each other, we must first define a set of hate subreddits to analyze that spans a variety of targeted identities. While researchers have previously curated exhaustive lists of subreddits that represent a particular ideology, such as the manosphere (Ribeiro et al. 2021) or the alt-right (Mamié, Horta Ribeiro, and West 2021), there currently exists no comprehensive dataset of hate subreddits that spans the diversity of ideologies that could be observed on the platform. For this reason, we create a large dataset of hate subreddits labeled by the identities they target. Our process consists of collecting a wide set of subreddits that are likely to be hate subreddits, detecting hate speech of different types in said subreddits, and clustering subreddits together that contain hate speech of similar types.
@@
To identify potential hate subreddits, we collect curated lists from various sources. First, we find lists of hate subreddits or banned subreddits previously identified in the literature (Mamié, Horta Ribeiro, and West 2021; Ribeiro et al. 2021; Vidgen et al. 2021a; Schmitz et al. 2024). As these papers turn to subreddits such as r/AgainstHateSubreddits to find potential hate groups, we supplement these lists with more subreddits by searching “list of hate subreddits” on r/AgainstHateSubreddits and finding each post listing hate subreddits within the first 100 search results as of October 2023. We then collect the entire history of comments and submissions from each subreddit using the Pushshift API (Baumgartner et al. 2020). All subreddits with fewer than 1000 users are removed from further analysis, as such subreddits represent less than 2% of all users in the dataset. Additionally, as many banned subreddits could be unrelated to hate (such as black market subreddits), one annotator (who is also an author of this paper) viewed 10 random comments and 10 random submissions from each subreddit, filtering out subreddits that were clearly unrelated to hate. At the end of this process, 168 subreddits remained.
@@
After collecting an initial set of potential hate subreddits, we define various types of hate speech within each subreddit based on the identities targeted by the speech, informed by “bias categories” in United States federal hate crime law  [^1] https://www.justice.gov/hatecrimes/learn-about-hate-crimes, where a hate crime is defined as “a crime motivated by bias against race, color, religion, national origin, sexual orientation, gender, gender identity, or disability.” We deviate from these bias categories by separating the “religion” category into two categories that represent speech directed at Jews and Muslims, as hate speech directed at these groups often additionally targets their race or national origin. We also add a category representing fat-shaming, as this type of hate speech is shown to have been prominent on Reddit (Chandrasekharan et al. 2017). Therefore, our final categories of hate speech are antisemitism, Islamophobia, ableism, fat-shaming, misogyny, xenophobia, racism, homophobia, and transphobia.
To find the prevalence of each type of hate speech in each subreddit, we utilize GPT-4, as it has been shown to perform well on the task of hate speech detection (Kumarage, Bhattacharjee, and Garland 2024). For each type of hate speech, we prompt GPT-4 to count the number of posts containing that type of hate speech in batches of 10 with the following string:
[LIST OF POSTS]
For each subreddit, we sample a total of 20 comments and 20 submissions to be submitted to GPT-4 for hate speech detection. The sample size was chosen as a trade-off between sufficiently sampling the subreddit, and avoiding excessive API costs as well as remaining within the input token limit of 128K tokens. Subreddits in which fewer than 10% of posts contain hate speech are removed from further analysis. This leaves us with a final set of 120 subreddits.
To validate our method of detecting hate speech, three annotators who are undergraduate or recently graduated students with a background in social science, annotated 10 comments and 10 submissions each from a set of 25 subreddits with which they were ethnographically familiar, having previously conducted research on the given subreddits. The annotators assessed the number of comments and submissions containing hate speech in each category for each subreddit, achieving a Krippendorff’s alpha score of 0.5 (Hayes and Krippendorff 2007), implying relatively poor inter-rater reliability. This low alpha score demonstrates the difficulty annotators have when assessing identities targeted, alike to the common issue of low annotator agreement in hatespeech classification (Vidgen et al. 2021b; Sachdeva et al. 2022). Comparing the average values of the human annotations to the estimates obtained by GPT-4, we achieve an overall $R^{2}$ value of 0.45 for our hate speech detection method. These results suggest GPT-4 is not perfect, but can still be a good proxy of human annotators.
To separate these subreddits into distinct categories, we employ $k$-means clustering. Each subreddit is assigned a vector representing the number of posts of each type of hate speech detected in the subreddit, normalized by its total number of posts containing hate speech. We run the $k$-means algorithm for values of $k$ ranging from two to 20 and choose the number of clusters that yield the highest silhouette score (three clusters). The average distributions of types of hate speech for each cluster are shown in Figure 2 and a UMAP plot of the distributions of hate for each subreddit is shown in Figure 1.
@@
We release the set of subreddits, their clusters, and the average number of posts containing hate speech of each category for each batch of posts, available at the following URL: https://github.com/dan-hickey1/peripatetic-hater. Upon publication, we intend to adhere to FAIR principles (FORCE11 2020) by making the data publicly available and easily accessible via GitHub. The data will be in a widely used format (CSV). Using the subreddit names, data from each subreddit can be obtained through Pushshift (Baumgartner et al. 2020). Caution should be exercised when using the dataset, as several subreddit names contain offensive phrases directed toward marginalized groups.
Figure 2: Distributions of the types of hate predicted by GPT-4 in each cluster.
@@
We label as peripatetic users those individuals who post in an alternative category of hate subreddit at some point after their initial post in a given hate subreddit. Because we seek to explore predictive markers of such peripatetic behavior after users have already participated in an antisocial environment, we restrict our focus to users who post in an alternative category of hate subreddit within six weeks of posting in their first hate subreddit, as we believe early peripatetic behavior will be more closely linked to the text expressed in the initial stages of hate subreddit participation. As the choice of this threshold is somewhat arbitrary, we compare individuals who become peripatetic users to users who do not subsequently post in additional categories of hate subreddits for all downstream analyses. For all analyses, we perform robustness checks varying this threshold between three weeks and one year.
@@
As automated accounts are likely to be ubiquitous across categories of subreddits but are not relevant to the questions at hand, we filter out probable bots from our dataset. Our method follows Schmitz et al. 2024 and is performed by removing accounts with the following keywords: “bot,” “auto,” “transcriber,” “gif,” “link,” and “twitter.” In total, our dataset contains approximately 3M users, 117K of whom post in an alternate category of hate subreddit within six weeks of their first post in a hate subreddit, and 2.5M of whom only post in one category of hate subreddit.

### v2

@@ Abstract
Many online hate groups exist to disparage others based on race, gender identity, sex, or other characteristics. The accessibility of these communities allows users to join multiple types of hate groups (e.g., a racist community and a misogynistic community), raising the question of whether users who join additional types of hate communities could be further radicalized compared to users who stay in one type of hate group. However, little is known about the dynamics of joining multiple types of hate groups, nor the effect of these groups on peripatetic users. We develop a new method to classify hate subreddits and the identities they disparage, then apply it to understand better how users come to join different types of hate subreddits. The hate classification technique utilizes human-validated deep learning models to extract the protected identities attacked, if any, across 168 subreddits. We find distinct clusters of subreddits targeting various identities, such as racist subreddits, xenophobic subreddits, and transphobic subreddits. We show that when users become active in their first hate subreddit, they have a high likelihood of becoming active in additional hate subreddits of a different category. We also find that users who join additional hate subreddits, especially those of a different category develop a wider hate group lexicon. These results then lead us to train a deep learning model that, as we demonstrate, usefully predicts the hate categories in which users will become active based on post text replied to and written. The accuracy of this model may be partly driven by peripatetic users often using the language of hate subreddits they eventually join. Overall, these results highlight the unique risks associated with hate communities on a social media platform, as discussion of alternative targets of hate may lead users to target more protected identities.
@@ Introduction
Recently, there has been extensive work documenting online hate communities and the impact that they have on the users of social media platforms (Schmitz et al. 2024; Russo et al. 2023). Such research is motivated by the combination of substantial increases in hate crimes (Federal Bureau of Investigation 2021), and the observation that radicalized individuals attribute their beliefs to online sources (Jensen et al. 2018). Much of the existing literature in this domain provides either case studies of individual communities (Chandrasekharan et al. 2017; Schmitz et al. 2024; Russo et al. 2023) or broad analyses of groups centered around a common theme, such as the manosphere (Ribeiro et al. 2021). Additionally, although prior work observes strong user overlap between anti-feminism and alt-right communities (Mamié, Horta Ribeiro, and West 2021), such research does not show how the core themes discussed in these communities are influenced by the users active in both communities.
Here, we extend the existing research corpus by curating a comprehensive dataset of 168 hate subreddits and using deep learning models and clustering techniques to group them by ideology (racist, misogynistic, xenophobic, transphobic, homophobic, Islamophobic, antisemitic, ableist, and general hate). In these subreddits, we analyze peripatetic users who participate in multiple categories of hate subreddits, attempting to characterize them and understand the impact that they have on Reddit’s ecosystem of hate. Our research questions are as follows:
@@ 4. RQ4
Our findings help answer each respective research question: (RQ1) Using matched-pair analysis, we find that participation in a given hate subreddit is associated with joining other hate subreddits at a rate roughly two times higher than that of matched users. (RQ2) Using BERTopic (Grootendorst 2022) and lexicons of ingroup language, we find that peripatetic users introduce into their initial subreddit topics of discussion from the categories in which they eventually participate. For example, a user who initially posts only in a misogynistic subreddit, but subsequently also posts in a racist subreddit, is more likely to introduce language characteristic of racist subreddits into the misogynistic subreddit. (RQ3) We find that users who join additional hate subreddits become more active in hate subreddits as a whole and use lexicons associated with these newly joined subreddits. Both results are more pronounced when users join hate subreddits of a category that differs from that of the first subreddit they joined. (RQ4) Finally, we develop a deep-learning model to predict the subreddit categories these peripatetic users will join. This task is challenging, as we are differentiating among users who already exhibit signs of hostility. We find that the language of users who eventually become peripatetic is strongly associated with the subreddit categories that they subsequently join.
Overall, our results offer a novel perspective on how hate communities may converge and become more “general” in an online environment as a result of peripatetic user activity. All data and code are in the following repository https://anonymous.4open.science/r/peripatetic-hater-1D26.
@@
We begin our exploration of online hate communities by building on other scholars’ prior research into specific categories of such communities; below we outline these categories and review work that has examined them. Misogynist. Extensive work has investigated the various identities targeted by online hate communities. One of the most widely studied groups is the manosphere, a collection of communities centered around misogyny and anti-feminism (Marwick and Caplan 2018; Habib, Srinivasan, and Nithyanand 2022). Despite sharing this common theme, there is a wide diversity of ideologies within the manosphere. Incels, for example, believe they are incapable of finding a sexual or romantic partner, and blame women and society for their misfortune (Daly and Reed 2022). This is in contrast with pick-up artists, who discuss strategies for seducing women, and “men going their own way” (MGTOW), who voluntarily decide to be celibate as they believe they should distance themselves from women (Lin 2017). A recent analysis shows MGTOW and incel communities are growing in popularity and are more toxic than other manosphere communities, which are decreasing in popularity (Ribeiro et al. 2021). Critically, members of incel communities have engaged in mass shooting events (Barcellona 2022).
Anti-LGBTQ. Anti-LGBTQ communities are also prevalent on the internet, including trans-exclusionary radical feminists (TERFs), an anti-transgender group. TERFs generally subscribe to a binary view of gender that is aligned with biological sex (disregarding intersex individuals), and see transgender people as a threat to women’s spaces (Jones and Slater 2020). Scholars have previously demonstrated flaws in such discourse and the harm it produces (Williams 2020). TERF communities have been present on Reddit in the past, with r/GenderCritical being the most prominent (Tiffany 2020). Previous work has found these communities to be more toxic than mainstream feminist communities (Balci et al. 2023), and models have been built to automatically identify the discourse employed by them (Lu and Jurgens 2022).
Racist/xenophobic. Racist groups have been highly prevalent on the internet, and on Reddit in particular (Chandrasekharan et al. 2017), as have a class of online groups known collectively as the alt-right, an extreme political movement that gained popularity through its utilization of the unique characteristics of social media (Winter 2019). The alt-right believes that the cultural identity of white individuals is at risk, with movements for social justice and political correctness being seen as principal threats (Center 2023). Alt-right online communities have radicalized users in part through the use of memes (Dafaure 2020), with one analysis showing that communities linked to the alt-right have a strong influence on the spread of memes in the larger web ecosystem (Zannettou et al. 2018). The misogynist movement “GamerGate” helped the alt-right movement gain adherents (Bezio 2018), demonstrating a relationship between seemingly distinct types of hate.
@@
Radicalization is explicable in terms of a number of sociological and psychological effects (Borum 2011; Maskaliūnaitė 2015; Neo 2016), which include: uncertainty reduction (Hogg, Kruglanski, and van den Bos 2013), achieving significance (Kruglanski et al. 2014), and the need to be popular (Siegel et al. 2011). A recent study theorized that many users initially participate in extreme political subreddits as they are disillusioned with a polarized political landscape and desire to hold beliefs outside of mainstream political viewpoints (Mann et al. 2023). Additionally, a study that found highly-upvoted posts on r/The_Donald were likely to contain anti-Muslim sentiment argued that the high number of upvotes normalized such viewpoints, strengthening the bonds of the ingroup as they united against an “other.” (Gaudette et al. 2021). A substantial corpus of existing research examines the micro-scale processes that drive people to radicalization. At the level of categories of online groups, scholars have studied whether certain communities are “gateways” to more extreme communities, such as the connection between the intellectual dark web and the alt-right (Ribeiro et al. 2020), or between the manosphere and the alt-right (Ribeiro et al. 2020; Barcellona 2022). Other studies measure traits indicative of user radicalization or extremism that are adopted by members of certain hate groups, finding that such traits increase once users participate in hate subreddits (Habib, Srinivasan, and Nithyanand 2022; Schmitz et al. 2024). Recommendation algorithms meanwhile may not necessarily contribute significantly to radicalization as users who consume extreme content on YouTube are driven to do so by external sources, and are characterized by high racial and gender resentment even before engaging with such content (Chen et al. 2023).
@@
A growing body of literature is devoted to understanding how users participate in multiple communities on the same platform. One theme in this research is the attempt to understand inter-community conflict, where members active in one community make posts in other communities in order to attack them (Efstratiou et al. 2023; Datta and Adar 2019). Some authors speculate that mass migrations of users among communities on Reddit may disrupt the linguistic evolution of destination communities, potentially influencing their likelihood of being banned (Habib et al. 2021). More generally, Rollo et al. (Rollo et al. 2022) devise a model of “attention flow” to quantitatively describe how users on Reddit shift their focus among subreddits over time. Scholars have also built models to predict whether users will participate in particular online communities (Jin, Lee, and Cheung 2010; Debaere, Coussement, and De Ruyck 2018), with, for example, one study finding that the social support that users in an online health community receive is predictive of their remaining in that community (Wang et al. 2017).
The present study builds on previous research by examining a multitude of hate communities while assessing the connections among all of them. In addition, we describe the tendency of users to subsequently join specific additional hate communities after initially joining a given hate community, and measure the user behavior that predates such additional participation. Finally, we build a model to predict which new hate subreddits users will join.
@@ Methods
To properly understand how different categories of hate subreddits influence each other, we define a set of hate subreddits that promulgate hate for a variety of identities. While researchers have previously curated exhaustive lists of subreddits that represent a particular ideology, such as the manosphere (Ribeiro et al. 2021) or the alt-right (Mamié, Horta Ribeiro, and West 2021), there currently exists no comprehensive dataset of hate subreddits that captures the diversity of ideologies that can be observed on the platform. We therefore create a large dataset of hate subreddits, each labeled by the identity or identities that it targets. Our process consists of first collecting a wide set of subreddits that are likely to be hate subreddits, detecting the targets of hate in each subreddit, and clustering subreddits together that contain hate speech of similar types.
@@
To identify potential hate subreddits, we collect curated lists from various sources. First, we find lists of hate subreddits or banned subreddits previously identified in the literature (Mamié, Horta Ribeiro, and West 2021; Ribeiro et al. 2021; Vidgen et al. 2021a; Schmitz et al. 2024). As these papers cite subreddits such as r/AgainstHateSubreddits to find potential hate groups, we supplement these lists with more subreddits by searching “list of hate subreddits” on r/AgainstHateSubreddits and finding each post listing hate subreddits within the first 100 search results as of October 2023. We then collect the entire history of comments and submissions from each potential hate subreddit using the Pushshift API (Baumgartner et al. 2020). All subreddits with fewer than 1,000 users are removed from further analysis in order to have sufficient data for each subreddit, this choice removes less than 2% of all users in the dataset. Additionally, as many banned subreddits could be unrelated to hate (such as subreddits concerned with the black market), one annotator (who is also an author of this paper) viewed 10 random comments and 10 random submissions from each subreddit, filtering out subreddits that were clearly unrelated to hate. At the end of this process, 168 subreddits remained. The links to all subreddits analyzed, as well as the finalized list of subreddits, are available in the repository.
@@
We next define identities targeted in each hate subreddit based on groups identified in United States federal hate crime law  [^1] https://www.justice.gov/hatecrimes/learn-about-hate-crimes, accessed 15 Sep, 2024, where a hate crime is defined as “a crime motivated by bias against race, color, religion, national origin, sexual orientation, gender, gender identity, or disability.” We then further divide the “religion” category into two categories, representing hate speech directed at Jews and hate speech directed at Muslims, respectively, as hate speech directed at these groups often additionally targets their race or national origin. Our final list of hate speech categories thus consists of antisemitism, Islamophobia, ableism, misogyny, xenophobia, racism, homophobia, and transphobia.
To find the prevalence of each category targeted in each subreddit, we train a deep learning model using the Measuring Hate Speech corpus (Sachdeva et al. 2022a). The Measuring Hate Speech corpus consists of 50k posts collected from X, YouTube, and Reddit. Each post has been coded by multiple annotators as to the identities targeted by the post, as well as whether the post constitutes hate speech (under several different levels/definitions of hate speech). To train our prediction model, we use the Demux architecture, a multi-label transformer-based architecture that generates embeddings for the names of labels, leveraging associations between related labels to improve prediction performance (Chochlakis et al. 2023). The names of the labels are prepended to each input sequence, and the text of the posts and labels are embedded using BERT-base. We note that after analyzing 50 posts per subreddit using the Google Translate API  [^2] https://cloud.google.com/translate, accessed 15 Sep, 2024, the average percentage of English posts per subreddit is 97%, implying an English-language BERT is appropriate.
These embeddings are input into a two-layer neural network with the sigmoid activation function applied to generate predictions. The model is fit using the binary cross-entropy loss function. We use a hidden layer size of 128, a dropout value of 0.1, and the Tanh activation function in the hidden layer based on similar hyperparameters that achieve high performance on this exact dataset (Sachdeva et al. 2022b), and we noticed minimal changes in performance when testing other values for the hyperparameters. As a baseline, we also train a basic BERT model with the same hyperparameters and compare the performance of the two models. We train both models using an NVIDIA RTX A6000 GPU.
Given that the “targeted identity” labels in the Measuring Hate Speech corpus are only intended to capture whether an identity group is mentioned by a post and not whether that post constitutes hate speech against the identity group (Sachdeva et al. 2022b), we combine a model to predict identity group label with other labels from the measuring hate speech corpus (negative sentiment, disrespect, insults, attacks on identity groups, and a canonical “hate speech” label). Based on the annotations described below, we found that the best performance was obtained using a model that defines hate against a targeted identity as one that detects both negative sentiment and a targeted identity in the text. All reports of model performance and analyses are therefore generated using this classifier. Other labels (i.e., disrespect, etc) yield similar results. To illustrate that these classifiers are similar, we display the matrix of Spearman correlations among the predicted probabilities of these classifiers in Table 5.
To validate our method of detecting hate speech, three human annotators, fluent in English, labeled 10 comments and 10 submissions each from a set of 25 random hate subreddits. The annotators assessed the number of comments and submissions containing hate speech in each category for each subreddit. To determine whether a given post constitutes hate speech, the annotators referred to the United Nations’ definition of hate speech [^3] https://www.un.org/en/hate-speech/understanding-hate-speech/what-is-hate-speech, accessed 15 Sep, 2024. The annotators achieved a Krippendorff’s alpha score of 0.52 (Hayes and Krippendorff 2007), implying relatively poor inter-rater reliability, consistent with other papers on hate speech (Vidgen et al. 2021b; Sachdeva et al. 2022a), and indicative of the difficulty to assess identities targeted. For each category, we calculate the $R^{2}$ value between the average values of human annotations and the number of posts predicted as hate speech for that category by Demux. Demux confidence thresholds for negative sentiment and identity targeted are varied to maximize the F1 score of predictions on the test set of the Measuring Hate Speech corpus. Table 3. To obtain estimates of uncertainty for the performance of our classifier, we train the model five separate times, using a different random split of 70% training data, 15% testing data, and 15% validation data each time. The thresholds that have the highest average F1 score across all five runs are chosen. When comparing predictions obtained via these thresholds to our human annotations, we obtain an average $R^{2}$ value of 0.49. The $R^{2}$ values of each category, as well as the values obtained using alternative hate speech detection approaches, are listed in Table 4.
In addition to using thresholds that maximize F1 scores in the Measuring Hate Speech corpus, we also consider the thresholds that maximize the average $R^{2}$ value of human annotations. This results in an average $R^{2}$ value of 0.77 across all categories. However, it is unknown whether the performance of these thresholds will generalize to the subreddits in our sample beyond the random 25 subreddits that we annotated, as the selected thresholds may be overfit to this set of subreddits. Given that it is unclear which set of thresholds will generalize to our entire set of subreddits, we repeat all downstream analyses using the set of thresholds that optimize the average $R^{2}$ value.
To separate these subreddits into distinct categories, we employ $k$-means clustering. For each subreddit, we sample and annotate 1K comments and 1K submissions. Each subreddit is assigned a vector of length eight, where each entry represents the proportion of posts targeting at a given identity. Because some categories, misogyny in particular, had a high false positive rate, we transform each estimate into a z-score where z0̄ means the proportion of posts targeting an identity is exactly the mean proportion of posts targeting that identity across all subreddits. We run the $k$-means algorithm for values of $k$ ranging from two to twenty and choose the number of clusters that yield the highest silhouette score (eight clusters). The z-scores for each cluster are shown in Figure 2 and a UMAP plot of the distributions of identities targeted for each subreddit is shown in Figure 1. If we vary the model thresholds to maximize the $R^{2}$ scores and cluster subreddits we find qualitatively similar clusters where the adjusted mutual information score between two sets of clusters is 0.62. We repeat all downstream analyses using this set of clusters and observe qualitatively similar results.
@@
We release the set of subreddits, their clusters, and the average number of posts containing hate speech of each category for each batch of posts, available at the following URL: https://anonymous.4open.science/r/peripatetic-hater-1D26. Upon publication, we intend to adhere to FAIR principles (FORCE11 2020) by making the data publicly available and easily accessible via GitHub. The data will be in a widely used format (CSV) or accessible via public repositories. Caution should be exercised when using the dataset, as several subreddit names contain offensive phrases directed toward marginalized groups.
Figure 2: Z-scores of the identities attacked in each cluster. Clusters are titled by the category of hate with the highest z-score (most over-represented).
@@
We label as peripatetic users those individuals who post in an alternative category of hate subreddit at some point after their initial post in a given hate subreddit. Because we seek to explore predictive markers of such peripatetic behavior after users have already participated in an antisocial environment, we restrict our focus to users who post in an alternative category of hate subreddit within six weeks of posting in their first hate subreddit, as we believe early peripatetic behavior will be more closely linked to the text expressed in the initial stages of hate subreddit participation. For robustness, we also show in the Appendix our results when setting the threshold to six months and when not setting any threshold.
@@
As automated accounts are likely to be ubiquitous across categories of subreddits but are not relevant to the questions at hand, we filter out probable bots from our dataset. Our method follows Schmitz et al. 2024 and is performed by removing accounts with the following keywords: “bot,” “auto,” “transcriber,” “gif,” “link,” and “twitter.” (Broadly, these qualitatively appear to remove many bot accounts. Organic accounts may include some of these words, such as Twitter, but we find the word “twitter” is so rarely used, $\sim 100$ accounts, as to not affect results either way). In total, our dataset contains approximately 4.2M users, 226K of whom post in an alternate category of hate subreddit within six weeks of their first post in a hate subreddit, and 3.3M of whom only post in one category of hate subreddit. We report more detailed statistics in Table 2.